The colorless solid phosphorus pentachloride reacts violently with water to give a mixture of phosphoric acid and hydrochloric acid .

(a) Write a balanced chemical equation for this reaction.

_____+______-------______+_______


(b) Determine the concentration (in moles per liter) of each of the acids that result from the complete reaction of 3.57×10-2 moles of phosphorus pentachloride with enough water to give a solution volume of 806 mL.

phosphoric acid : __________ M.

hydrochloric acid : ___________ M.
1
Expert's answer
2019-03-12T15:54:36-0400

PCl5 + 4H2O = H3PO4 + 5HCl

n(H3PO4) = 3.57*10-2 mol

n(HCl) = 5*3.57*10-2 mol = 0.1785 mol

C(H3PO4) = n/V = 3.57*10-2 mol / 0.806 L = 0.0443 mol/L

C(HCl) = 0.1785 mol / 0.806 L = 0.2215 mol/L

phosphoric acid: 0.0443 M

hydrochloric acid: 0.2215 M